OpenStudy (anonymous):

Input the equation for the following line in standard form Ax + By = C. Reduce all fractional answers to lowest terms.A line that passes through the origin and is parallel to x + y = 6. Input the equation for the following line in standard form Ax + By = C. Reduce all fractional answers to lowest terms.A line that passes through the origin and is parallel to x + y = 6. @Mathematics

OpenStudy (saifoo.khan):

Parallel means same slope. y = -x + 6 slope = -1.. y = -x

OpenStudy (anonymous):

that's it?

OpenStudy (saifoo.khan):

Yes.

OpenStudy (anonymous):

Well it is wrong.. I take some online courses and it says that I am wrong.

OpenStudy (saifoo.khan):

wait, the answer is correct, but the form is not. y = -x x + y = 0 <--- insert that.

OpenStudy (anonymous):

Bravo!

OpenStudy (saifoo.khan):

=D

OpenStudy (anonymous):

Can you do another Por favor?

OpenStudy (saifoo.khan):

Sure.

OpenStudy (anonymous):

Input the equation for the following line in standard form Ax + By = C. Reduce all fractional answers to lowest terms.A line with y -intercept of 2 that is parallel to 2 x + y = -5.

OpenStudy (anonymous):

Step by step please...

OpenStudy (saifoo.khan):

2 x + y = -5. Parallel to that line. this means that the slope must be equal.. 2 x + y = -5. y = -2x -5 Slope = -2. Now, y = mx + b y = -2x + 2

OpenStudy (anonymous):

Ok so 2x-2=.... I am feeling really dumb.

OpenStudy (saifoo.khan):

Yes, now it wants in standard form, so: y = -2x + 2 2x + y = 2.

OpenStudy (anonymous):

Oh, that makes sense. Ok here is the last one, then I will try some myself. Input the equation for the following line. Reduce all fractional answers to lowest terms.A line that passes midway between and is parallel to both x - y = 6 and x - y = -6.

OpenStudy (anonymous):

This one seems different.

OpenStudy (saifoo.khan):

Yes it is.. It says in b/w. if u see, the y intercepts are -6 and 6. the midpoint is 0 b/w them. it means that it passes through the origin. and since they are parallel, so same slope. slope = 1. y = x y - x = 0

OpenStudy (anonymous):

It's partially correct, but I got 3 out of 5.

OpenStudy (saifoo.khan):

in this one?

OpenStudy (saifoo.khan):

or in total?

OpenStudy (anonymous):

Yeah

OpenStudy (saifoo.khan):

Why is that? o.O

OpenStudy (anonymous):

No Idea.

OpenStudy (saifoo.khan):

Did u wanted full step by step working?

OpenStudy (saifoo.khan):

Coz mostly online people i teach, they dont want it.

OpenStudy (anonymous):

Sure that would make sense... I could show it to my teacher tomorrow.

OpenStudy (anonymous):

Well I am different. Lol

OpenStudy (saifoo.khan):

LOL, Yea.. right.

OpenStudy (anonymous):

Hello...

OpenStudy (saifoo.khan):

Helloo..?

OpenStudy (anonymous):

Yeah were you able to do the math? I was trying to figure it out. Lol

OpenStudy (saifoo.khan):

Should i do it again?

OpenStudy (anonymous):

Yeah that would be helpful!

OpenStudy (saifoo.khan):

if u see, the y intercepts are -6 and 6. Distance = 12. 12/2 = 6. -6+6 = 0 the midpoint is 0 b/w them. it means that it passes through the origin. Origin = (0,0) and since they are parallel, so same slope. slope = 1. y = x + 0 y - x = 0